A rectangular prism has sides that are in the ratio of 3:5:9. If the volume of the prism is 8640 cm2, then the surface area of the prism, in square millimetres, is...
Answer by CubeyThePenguin(3113)

(3x)(5x)(9x) = 8640
135x^3 = 8640
x^3 = 64

x = 4
sides: 12, 20, 36

surface area: 2(12 * 20 + 12 * 36 + 20 * 36) = 2,784 cm^2 = 278,400 mm^2